A ladder that is 16 ft long is leaning against the side of a building. If the angle formed between the ladder and the ground is 72 degrees, how far is the bottom of the ladder from the base of the building?


1
Expert's answer
2022-02-21T15:25:51-0500

In the situation described above, a right-angle triangle is formed where, the length of the ladder forms the hypotenuse "H=16" , from the base of the building to the bottom of the ladder is the base and from the bottom of the building to the ladder's top is the height.

The angle formed between the ladder and the ground is "72\degree". The distance to the bottom of the ladder from the base of the building is the base (B) given as, "B=Cos(72\degree)\times H=0.3090\times 16=4.944" ft.